The viscosity between two parallel plates with a distance of 1 mm between them is 0.01 kg/mxs is an oil. Bottom plate 3 m/S to the left base plate right (2+A) M/S is rapidly pulled. a) find the force that should be applied to the top plate with a field of 1+B(m^2) as newton. b ) find the location ( distance from the lower surface) of the point where the fluid velocity is.chegg
1
Expert's answer
2020-12-22T05:04:19-0500

(a) Viscosity between two plates is 0.01 Ns/m2, distance between plate is = 1mm= 0.001 m

velocity with which plate pulled is = 3m/s , Area of the plate is (1+B) m2

we know that in case viscous force acting between plates

we have formula as

"F= A \mu \frac{du}{dy}"


"F= (1+B)\times 0.01\times \frac{3}{0.001}=30 (1+B) N"


Force needed to pull plates= 30 (1+B) N


(b) In b part question is not complete.